# How Do Document Parsers Break? Auditing Structural Vulnerability in Document Intelligence

### 19 May 2026 v1

@@ B.6 B-SLR Decomposition Algorithm
Algorithm 1 gives the implementation-level procedure used to decompose B-SLR into direct-occlusion and topology-dominant components. Within the topology-dominant subset $\mathcal{U}_{\mathrm{topo}}$, failures are further assigned to three mutually exclusive categories under a fixed priority order: merge, misclass, and degraded. Merge captures multiple clean elements assigned to the same perturbed prediction; misclass captures retained spatial correspondence with a changed canonical category; and degraded covers the remaining non-occluded failures, including text degradation without direct physical coverage.
Input: $E,\;E_{\mathrm{adv}},\;S(P)$
@@ 6 $\hat{e}\leftarrow m(e)$
8 if $\hat{\iota}<\tau_{\mathrm{iou}}$ then
9    if $\rho(e,S(P))\geq\eta_{\mathrm{occ}}$ then miss
10    else degraded
11
@@
17
18 else if $\operatorname{txt}(e)\neq\epsilon$ and $\mathrm{TextSim}(e,\hat{e})<\tau_{\mathrm{text}}$ then
19    if $\rho(e,S(P))\geq\eta_{\mathrm{occ}}$ then miss
20    else degraded
21
22 else intact
23
@@ C.1 Auditing Protocol Overview and Identifier Convention
where $\mathsf{P}_{n}$ identifies the probe family in Table 6, $t\in\{\mathtt{a},\mathtt{c},\mathtt{r},\mathtt{b}\}$ denotes placement, and $\xi$ records the dominant varied parameter. Tables 11, 12, and 13 provide the full decoding.
@@ C.2 Phase 1 Fixed Configurations
A07 $\mathsf{P}_{4}\!\cdot\!\mathtt{c}[5\%]$ $a\!=\!5\%,\;\beta\!=\!0.3$ erasure mild
A08 $\mathsf{P}_{4}\!\cdot\!\mathtt{c}[20\%]$ $a\!=\!20\%,\;\beta\!=\!1.0$ erasure severe
@@
A19 $\mathsf{P}_{4}\!\cdot\!\mathtt{b}[20\%]$ $a\!=\!20\%,\;\beta\!=\!1.0$ erasure bridge
@@ C.3 EIR-Targeted NT Configurations
Table 12: EIR-targeted configurations under the NT01–NT07 identifier convention. The probe family and placement mechanism are fixed; only the target interference ratio $\tau$ varies.
@@ C.4 Randomized Sweep Configurations
S04 P4 $a\!\sim\!U(3\%,25\%),\;\beta\!\sim\!U(0.2,1)$ content
@@ C.5 Inference, Randomness, and Software
Backend DeepSeek-Chat Qwen-VL-Max / GPT-4o
@@ Appendix D Additional Results and Robustness Checks
D.1 Channel-by-Gradient Breakdown
Table 16 expands the gradient analysis in Section 5.2.4 by reporting configuration-level $R^{2}$ between each diagnostic descriptor and each B-SLR channel. This breakdown complements Table 4 by showing how the five exposure descriptors relate to composite B-SLR and its IoU-fail and text-fail-only subchannels.
@@ D.1 Channel-by-Gradient Breakdown
Table 16: Channel-by-gradient regression breakdown: configuration-level $R^{2}$ from univariate OLS fits between each diagnostic descriptor and each B-SLR channel. Bold marks the strongest predictor per row.
@@ E.1 Reproducibility Checklist
  • Configurations: A01–A22 and NT01–NT07 for Phase 1 fixed auditing (Tables 11 and 12); S01–S13 for randomized sweeps (Table 13); five policy families for Phase 2.
@@ E.2 Prompt Templates
The policy prompts are designed to emit the same output schema: a probe type, bounded probe parameters, and a placement strategy. The biased and neutral LLM variants differ in role framing, objective wording, strategy labels, and context rendering, as summarized in Table 7. Figure 7 shows abridged prompt excerpts for the three prompt-based policy variants. For the text-only policies, the biased variant uses a structure-aware page description through context.to_text_description(), whereas the neutral variant uses a coordinate-only rendering through context.to_neutral_description(). The VLM policy instead receives the raw page image and predicts only the probe type and placement strategy.
(a) LLM-biased prompt excerpt You are an expert adversarial tester for Document Layout Analysis (DLA) systems.
Your goal is to choose ONE visual probe and place it on the document image to cause maximum parsing failure, such as merging, splitting, or missing layout blocks.
Available probe types: P1--P9.
Placement strategies: bridge / anchor / content / random.
Output strictly valid JSON.
Context: structure-aware page description with block coordinates, gap information, and candidate vulnerable regions.
(b) LLM-neutral prompt excerpt You are a quality assurance analyst evaluating the robustness of Document Layout Analysis (DLA) systems.
Select ONE visual perturbation probe and one placement strategy to test how the model responds to realistic document degradation.
Placement options: between / edge / inside / anywhere.
Output strictly valid JSON.
Context: coordinate-only page description with block types and bounding boxes, without explicit structural vulnerability hints.
(c) VLM strategy-only prompt excerpt Look at this document image carefully. Observe the text blocks, figures, tables, and whitespace regions.
Based only on what you see in the image, select a probe type and a placement strategy.
Placement options: between / edge / inside / anywhere.
The system will automatically determine exact placement coordinates based on the chosen strategy.
Output strictly valid JSON.
Input: raw page image only; no coordinate text description is provided.
@@ E.3 Context Encoder for Non-VLM Policies
The biased prompt uses a structure-aware text rendering through to_text_description(), while the neutral prompt uses a coordinate-only rendering through to_neutral_description().

### 26 May 2026 v2

@@ B.6 B-SLR Decomposition Algorithm
Algorithm 1 gives the implementation-level procedure used to decompose B-SLR into direct-occlusion and topology-dominant components. The algorithm first identifies structurally failed elements using the same IoU–TextSim gates as the main B-SLR definition. Failed elements are then split into direct-occlusion and topology-dominant subsets according to the coverage ratio $\rho(e,P)$. Within the topology-dominant subset $\mathcal{U}_{\mathrm{topo}}$, failures are further assigned to three mutually exclusive categories under a fixed priority order: merge, misclass, and degraded. Merge captures multiple clean elements assigned to the same perturbed prediction; misclass captures retained spatial correspondence with a changed canonical category; and degraded covers the remaining non-occluded failures, including text degradation without direct physical coverage.
Input: $E,\;E_{\mathrm{adv}},\;A(P)$
@@ 6 $\hat{e}\leftarrow m(e)$
[…]
9 $\mathrm{failed}\leftarrow(\mathrm{iou}<\tau_{\mathrm{iou}})\vee(\hat{s}<\tau_{\mathrm{text}})$
10
11  1exif not $\mathrm{failed}$ then
12    intact
13
14 else if $\rho(e,P)\geq\eta_{\mathrm{occ}}$ then
15    miss
16
@@ 9 $\mathrm{failed}\leftarrow(\mathrm{iou}<\tau_{\mathrm{iou}})\vee(\hat{s}<\tau_{\mathrm{text}})$
22
23 else
24    degraded
25
26   end if
27
@@
B.7 Policy-Audit Metric Shorthands
[…]
$\mathrm{TopoShare}=\frac{\mathrm{SLR}_{\mathrm{topo}}}{\mathrm{SLR}_{\mathrm{miss}}+\mathrm{SLR}_{\mathrm{topo}}}.$ (40)
Eff-B and Eff-C denote per-footprint efficiency for structural loss and OCR instability:
$\mathrm{Eff\mbox{-}B}=\frac{\mathrm{B\mbox{-}SLR}}{\mathrm{TOR}},\qquad\mathrm{Eff\mbox{-}C}=\frac{\mathrm{CER}}{\mathrm{TOR}}.$ (41)
These efficiency scores are used only as relative indicators under comparable small-footprint regimes; they should be interpreted together with absolute damage and pathway composition.
@@ C.1 Auditing Protocol Overview and Identifier Convention
where $\mathsf{P}_{n}$ identifies the probe family in Table 5, $t\in\{\mathtt{a},\mathtt{c},\mathtt{r},\mathtt{b}\}$ denotes placement, and $\xi$ records the dominant varied parameter. Tables 10, 11, and 12 provide the full decoding.
@@ C.2 Phase 1 Fixed Configurations
A07 $\mathsf{P}_{4}\!\cdot\!\mathtt{c}[5\%]$ $a_{\mathrm{area}}\!=\!5\%,\;\beta\!=\!0.3$ erasure mild
A08 $\mathsf{P}_{4}\!\cdot\!\mathtt{c}[20\%]$ $a_{\mathrm{area}}\!=\!20\%,\;\beta\!=\!1.0$ erasure severe
@@
A19 $\mathsf{P}_{4}\!\cdot\!\mathtt{b}[20\%]$ $a_{\mathrm{area}}\!=\!20\%,\;\beta\!=\!1.0$ erasure bridge
@@ C.3 EIR-Targeted NT Configurations
Table 11: EIR-targeted configurations under the NT01–NT07 identifier convention. The probe family and placement mechanism are fixed; only the target $\tau$ varies.
@@ C.4 Randomized Sweep Configurations
S04 P4 $a_{\mathrm{area}}\!\sim\!U(3\%,25\%),\;\beta\!\sim\!U(0.2,1)$ content
@@ C.5 Inference, Randomness, and Software
Backend DeepSeek-Chat Qwen-VL-Max
@@ Appendix D Additional Results and Robustness Checks
D.1 Channel-by-Exposure Breakdown
Table 15 expands the exposure analysis in Section 5.2.4 by reporting configuration-level $R^{2}$ between each diagnostic descriptor and each B-SLR channel. This breakdown complements Table 3 by showing how the five exposure descriptors relate to composite B-SLR and its IoU-fail and text-fail-only subchannels.
@@ D.1 Channel-by-Exposure Breakdown
Table 15: Channel-by-exposure regression breakdown: configuration-level $R^{2}$ from univariate OLS fits between each diagnostic descriptor and each B-SLR channel. Bold marks the strongest predictor per row.
@@ E.1 Reproducibility Checklist
  • Configurations: A01–A22 and NT01–NT07 for Phase 1 fixed auditing (Tables 10 and 11); S01–S13 for randomized sweeps (Table 12); five policy families for Phase 2.
@@ E.2 Prompt Templates
LLM-biased prompt excerpt. Role: expert adversarial tester for DLA. Goal: choose one visual probe and placement to maximize parsing failure, including merge, split, or missing-block errors. Probes: P1–P9; strategies: bridge/anchor/content/random; output: valid JSON. Context: structure-aware page description with block coordinates, gap information, and candidate vulnerable regions.
LLM-neutral prompt excerpt. Role: quality-assurance analyst for DLA robustness. Goal: select one realistic perturbation probe and placement strategy to test model response. Strategies: between/edge/inside/anywhere; output: valid JSON. Context: coordinate-only page description with block types and bounding boxes, without explicit structural vulnerability hints.
VLM strategy-only prompt excerpt. Input: raw page image only. Goal: inspect visible text blocks, figures, tables, and whitespace, then select a probe type and placement strategy. Exact coordinates are determined automatically from the chosen strategy; output: valid JSON.